OpenStudy (anonymous):

Solve and check answer 9/(2x+1)=4/x

OpenStudy (anonymous):

\[\frac{ 9 }{2x+1 } = \frac{ 4 }{ x }\] cross multiplying 9x = 4(2x+1) solve for x